OpenStudy (anonymous):

Statement: m<3+m<2=90 Reason?

OpenStudy (mathstudent55):

What statements came before that?

OpenStudy (anonymous):

<3 AND <2 ARE COMPLEMENTARY. REASON: GIVEN M<1+m<2 = 90 REASON:GIVEN

OpenStudy (anonymous):

Given <3 and <2 are complementary. m<1+m<2=90. Prove <3 not equal to <1

Nnesha (nnesha):

what does complementary mean do you know that ??

OpenStudy (anonymous):

two angles whose measures have the sum of 90

OpenStudy (anonymous):

no angle 2 and 3 have to have a sum of 90 to be complementary\

Nnesha (nnesha):

do you have any diagram with it ??

OpenStudy (mathstudent55):

Since angles 3 and 2 are complementary, then by the definition of complementary angles, you can state that m<3 + m<2 = 90.

OpenStudy (mathstudent55):

Statements | Reasons ------------------------------------|----------------------------------- 1. <3 and <2 are complementary | 1. Given 2. m<1+ m<2 = 90 | 2. Given 3. m<3 + m<2 = 90 | 3. Definition of complementary angles